### v1

# FALCON: An ML Framework for Fully Automated Layout-Constrained Analog Circuit Design

Abstract
Designing analog circuits from performance specifications is a complex, multi-stage process encompassing topology selection, parameter inference, and layout feasibility. We introduce FALCON, a unified machine learning framework that enables fully automated, specification-driven analog circuit synthesis through topology selection and layout-constrained optimization. Given a target performance, FALCON first selects an appropriate circuit topology using a performance-driven classifier guided by human design heuristics. Next, it employs a custom, edge-centric graph neural network trained to map circuit topology and parameters to performance, enabling gradient-based parameter inference through the learned forward model. This inference is guided by a differentiable layout cost, derived from analytical equations capturing parasitic and frequency-dependent effects, and constrained by design rules. We train and evaluate FALCON on a large-scale custom dataset of 1M analog mm-wave circuits, generated and simulated using Cadence Spectre across 20 expert-designed topologies. Through this evaluation, FALCON demonstrates >99% accuracy in topology inference, <10% relative error in performance prediction, and efficient layout-aware design that completes in under 1 second per instance. Together, these results position FALCON as a practical and extensible foundation model for end-to-end analog circuit design automation. Our code and dataset are publicly available at https://github.com/AsalMehradfar/FALCON.
1 Introduction
Analog radio frequency (RF) and millimeter-wave (mm-wave) circuits are essential to modern electronics, powering critical applications in signal processing [1], wireless communication [2], sensing [3], radar [4], and wireless power transfer systems [5]. Despite their importance, the design of analog circuits remains largely manual, iterative, and dependent on expert heuristics [6, 7, 8]. This inefficiency stems from several challenges: a vast and continuous design space that is difficult to explore systematically; tightly coupled performance metrics (e.g. gain, noise, bandwidth, and power) that create complex trade-offs; and physical and layout-dependent interactions that complicate design decisions. As demand grows for customized, high-performance analog blocks, this slow, expert-driven design cycle has become a critical bottleneck. While machine learning (ML) has revolutionized digital design automation, analog and RF circuits still lack scalable frameworks for automating the full pipeline from specification to layout.
While recent ML approaches have made progress in analog circuit design, they typically target isolated sub-tasks such as topology generation or component sizing [9, 10] at the schematic level, without addressing the full synthesis pipeline. Many efforts assume fixed topologies [11, 12, 13, 14], limiting adaptability to new specifications or circuit families. Optimization strategies often rely on black-box methods that do not scale well to large, continuous design spaces [15]. Some methods predict performance metrics directly from netlists [16], but do not support inverse design, i.e., generating circuit parameters from target specifications. Furthermore, layout awareness is typically handled as a separate post-processing step [17], missing the opportunity to guide optimization with layout constraints. Finally, many available benchmarks are built on symbolic or synthetic simulations [18], lacking the fidelity and realism of the process of commercial grade design flows. As a result, current ML pipelines do not allow fully generalizable, layout-aware, and end-to-end analog circuit design.
We propose FALCON (Fully Automated Layout-Constrained analOg circuit desigN), a scalable and modular machine learning framework for end-to-end analog and RF circuit design. Built on a dataset of over one million Cadence-simulated circuits, FALCON comprises three core components (Figure 1): (1) a lightweight multilayer perceptron (MLP) selects the most appropriate topology given a target performance specification; (2) a generalizable graph neural network (GNN) maps circuit topology and element-level parameters to performance metrics, operating on a native graph representation derived from Cadence netlists; and (3) gradient-based optimization over the forward GNN model recovers design parameters that meet the target specification, guided by a differentiable layout-aware loss that encodes parasitic effects and physical constraints. Notably, the GNN model in FALCON generalizes effectively to unseen topologies, enabling inverse design across diverse circuit families, even in low-data regimes, with optional fine-tuning for improved accuracy. By integrating layout modeling directly into the optimization process, FALCON unifies schematic and physical considerations within a single differentiable learning framework.
Our main contributions are as follows:
  • We construct a large-scale analog/RF circuit dataset comprising over one million Cadence-simulated datapoints across 20 expert-designed topologies and five circuit types.
  • We introduce a native netlist-to-graph representation that preserves both structural and parametric fidelity, enabling accurate learning over physical circuit topologies.
  • We develop a modular ML framework for end-to-end inverse design, incorporating performance-driven topology selection and layout-aware gradient-based optimization, with a differentiable loss that enforces area constraints, design-rule compliance, and frequency-dependent modeling of passive components.
  • We design a generalizable GNN capable of accurate performance prediction and parameter inference across both seen and unseen topologies, with optional fine-tuning.
Figure 1: Our AI-based circuit design pipeline. Given a target performance specification, FALCON first selects a suitable topology, then generates design parameters through layout-aware gradient-based reasoning with GNN model. Then, the synthesized circuit is validated using Cadence simulations.
2 Related Work
While recent ML-based approaches have advanced analog and RF circuit design, they typically target isolated stages of the design flow—such as topology generation, parameter sizing, or schematic-level performance prediction—without supporting unified, end-to-end synthesis. FALCON bridges this gap by jointly addressing aforementioned stages within a single framework.
Topology generation methods aim to select or synthesize candidate circuit structures [9, 19, 20], often using discrete optimization or generative models to explore the circuit graph space. However, these approaches typically target low-frequency or simplified designs [9] and may produce physically invalid or non-manufacturable topologies. In contrast, FALCON leverages a curated set of netlists, ensuring manufacturable validity and eliminating the need to rediscover fundamental circuit structures.
Parameter sizing and performance prediction have been explored through various learning paradigms. Reinforcement learning [10, 21] and Bayesian optimization [15, 22] optimize parameters via trial-and-error, often requiring large simulation budgets. Supervised learning methods [23, 24, 11] regress parameter values from performance targets under fixed topologies. Graph-based models [16] incorporate topology-aware representations to predict performance metrics from netlists. However, these approaches focus on forward prediction or black-box sizing and do not support inverse design across varied topologies. In contrast, FALCON unifies forward modeling and parameter inference in a single differentiable architecture that generalizes to unseen netlists.
Layout-aware sizing and parasitic modeling have been explored to mitigate schematic-to-layout mismatch. Parasitic-aware methods [25] integrate pre-trained parasitic estimators into Bayesian optimization loops for fixed schematics. While effective for estimation, these approaches rely on time-consuming black-box search and lack inverse design capabilities. Other methods, such as ALIGN [26] and LayoutCopilot [27], generate layouts from fully sized netlists using ML-based constraint extraction or scripted interactions, but assume fixed parameters and do not support co-optimization or differentiable inverse design. In contrast, FALCON embeds layout objectives directly into the learning loss, enabling joint optimization of sizing and layout without relying on external parasitic models. For mm-wave circuits, our layout-aware loss captures frequency-sensitive constraints via simplified models that implicitly reflect DRC rules, EM coupling, and performance-critical factors such as quality factor and self-resonance frequency.
Datasets for analog design are often limited to symbolic SPICE simulations or small-scale testbeds that do not reflect real-world design flows. AnalogGym [18] and AutoCkt [13] rely on synthetic circuits and symbolic simulators, lacking the process fidelity, noise characteristics, and layout-dependent behavior of foundry-calibrated flows. In contrast, FALCON is trained on a large-scale dataset constructed from over one million Cadence-simulated circuits across 20 topologies and five circuit categories, offering a substantially more realistic foundation for ML-driven analog design.
To the best of our knowledge, FALCON is the first framework to unify topology selection, parameter inference, and layout-aware optimization in a single end-to-end pipeline, validated at scale using industrial-grade Cadence simulations for mm-wave analog circuits.
3 A Large-Scale Dataset and Inverse Design Problem Formulation
3.1 Dataset Overview
We construct a large-scale dataset of analog and RF circuits simulated using industry-grade Cadence tools [28] with a 45nm CMOS process design kit (PDK). The dataset spans five widely used mm-wave circuit types for wireless applications [29, 30]: low-noise amplifiers (LNAs) [31, 32, 33, 34], mixers [35, 36, 37, 38], power amplifiers (PAs) [39, 40, 41, 42, 43], voltage amplifiers (VAs) [44, 45, 46, 47, 48], and voltage-controlled oscillators (VCOs) [49, 50, 51, 52, 53]. Each circuit type is instantiated in four distinct topologies, resulting in a total of 20 expert-designed architectures.
For each topology, expert-designed schematics were implemented in Cadence Virtuoso, and key design parameters were manually identified based on their functional relevance. Parameter ranges were specified by domain experts and systematically swept using Cadence ADE XL, enabling parallelized Spectre simulations across the design space. For each configuration, performance metrics—such as gain, bandwidth, and oscillation frequency—were extracted and recorded. Each datapoint therefore includes the full parameter vector, the corresponding Cadence netlist, and the simulated performance metrics. The resulting dataset comprises over one million datapoints, capturing a wide range of circuit behaviors and design trade-offs across diverse topologies. This large-scale, high-fidelity dataset forms the foundation for training and evaluating our inverse design pipeline.
3.2 Graph-Based Circuit Representation
To enable flexible and topology-agnostic learning, we represent each analog circuit as a graph extracted from its corresponding Cadence netlist. Nodes correspond to voltage nets (i.e., electrical connection points), and edges represent circuit elements such as transistors, resistors, capacitors, or sources. Multi-terminal devices—such as transistors and baluns—are decomposed into multiple edges, and multiple components may connect the same node pair, resulting in heterogeneous, multi-edged graphs that preserve structural and functional diversity.
Recent works such as DICE [54] have explored transistor-level circuit-to-graph conversions for self-supervised learning, highlighting the challenges of faithfully capturing device structure and connectivity. In contrast, our approach maintains a native representation aligned with foundry-compatible netlists. Rather than flattening or reinterpreting device abstractions, we preserve symbolic parameters, multi-edge connections, and device-specific edge decomposition directly from the schematic source, enabling scalable learning across diverse analog circuit families.
To support learning over such structured graphs, each edge is annotated with a rich set of attributes: (i) a categorical device type, specifying the component and connected terminal pair (e.g., NMOS drain–gate, resistor); (ii) numeric attributes, such as channel length or port resistance, fixed by the schematic; (iii) parametric attributes, defined symbolically in the netlist (e.g., W1, R3) and resolved numerically during preprocessing; (iv) one-hot categorical features, such as source type (DC, AC, or none); and (v) computational attributes, such as diffusion areas (Ad, As) derived from sizing. This rule-based graph construction generalizes across circuit families without task-specific customization. Graphs in the FALCON dataset range from 4–40 nodes and 7–70 edges, reflecting the variability of practical analog designs.
3.3 Inverse Design Problem Definition
In analog and RF circuit design, the traditional modeling process involves selecting a topology $T$ and parameter vector $x$, then evaluating circuit behavior via simulation to obtain performance metrics $y=f(T,x)$. This forward workflow depends heavily on designer intuition, manual tuning, and exhaustive parameter sweeps. Engineers typically simulate many candidate $(T,x)$ pairs and select the one that best satisfies the target specification—a slow, costly, and unguided process.
In contrast, our goal is to perform inverse design: given a target performance specification $y_{\text{target}}$, we aim to directly infer a topology and parameter configuration $(T,x)$ such that $f(T,x)\approx y_{\text{target}}$, without enumerating the full design space. This inverse problem is ill-posed and the search space is constrained by both device-level rules and layout-aware objectives.
Formally, the task is to find the optimal topology $T^{*}\in\mathcal{T}$ and the optimal parameters $x^{*}\in\mathbb{R}^{p}$ such that $f(T^{*},x^{*})\approx y_{\text{target}}$ where $f:\mathcal{T}\times\mathbb{R}^{p}\rightarrow\mathbb{R}^{d}$ the true performance function implemented by expensive Cadence simulations. In practice, $f$ is nonlinear and non-invertible, making direct inversion intractable. FALCON addresses this challenge through a modular, three-stage pipeline:
Stage 1: Topology Selection. We frame topology selection as a classification problem over a curated set of $K$ candidate topologies $\{T_{1},\dots,T_{K}\}$. Given a target specification $y_{\text{target}}$, a lightweight MLP selects the topology $T^{*}\in\mathcal{T}$ most likely to satisfy it, reducing the need for exhaustive search.
Stage 2: Performance Prediction. Given a topology $T$ and parameter vector $x$, we train a GNN $f_{\theta}$ to predict the corresponding performance $\hat{y}=f_{\theta}(T,x)$. This model emulates the forward behavior of the simulator $f$, learning a continuous approximation of circuit performance across both seen and unseen topologies. By capturing the topology-conditioned mapping from parameters to performance, $f_{\theta}$ serves as a differentiable surrogate that enables gradient-based inference in the next stage.
Stage 3: Layout-Aware Gradient Reasoning. Given $y_{\text{target}}$ and a selected topology $T^{*}$, we infer a parameter vector $x^{*}$ by minimizing a loss over the learned forward model $f_{\theta}$. Specifically, we solve:
$x^{*}=\arg\min_{x}\;\mathcal{L}_{\text{perf}}(f_{\theta}(T^{*},x),y_{\text{target}})+\lambda\,\mathcal{L}_{\text{layout}}(x),$ (1)
where $\mathcal{L}_{\text{perf}}$ measures prediction error, and $\mathcal{L}_{\text{layout}}$ encodes differentiable layout-related constraints such as estimated area and soft design-rule penalties. Optimization is performed via gradient descent, allowing layout constraints to guide the search through a physically realistic parameter space.
4 Stage 1: Performance-Driven Topology Selection
Task Setup. We formulate topology selection as a supervised classification task over a fixed library of 20 expert-designed circuit topologies $\mathcal{T}=\{T_{1},T_{2},\dots,T_{20}\}$. Rather than generating netlists from scratch—which often leads to invalid or impractical circuits—we select from a vetted set of designer-verified topologies. This ensures that all candidates are functionally correct, layout-feasible, and manufacturable. While expanding the topology set requires retraining, our lightweight MLP classifier enables rapid updates, making the approach scalable. This formulation also aligns with practical design workflows, where quickly identifying a viable initial topology is critical.
Each datapoint is represented by a 16-dimensional performance vector of key analog/RF metrics [^1] DC power consumption (DCP), voltage gain (VGain), power gain (PGain), conversion gain (CGain), $\text{S}_{11}$, $\text{S}_{22}$, noise figure (NF), bandwidth (BW), oscillation frequency (OscF), tuning range (TR), output power (OutP), $\text{P}_{\text{SAT}}$, drain efficiency (DE), power-added efficiency (PAE), phase noise (PN), voltage swing (VSwg).. We normalize features using z-scores computed from the training set. Missing metrics (e.g., oscillation frequency for amplifiers) are imputed with zeros, yielding zero-centered, fixed-length vectors that retain task-relevant variation. Dataset splits are stratified to preserve class balance across training, validation, and test sets. We assume each target vector is realizable by at least one topology in $\mathcal{T}$, though the library can be extended with new designs.
Model Architecture and Training. We train a 5-layer MLP with hidden size 256 and ReLU activations for this problem. The model takes the normalized performance vector $y_{\text{target}}\in\mathbb{R}^{16}$ as input and outputs a probability distribution over 20 candidate topologies. The predicted topology is selected as $T^{*}=\arg\max_{T_{k}\in\mathcal{T}}\ \text{MLP}(y_{\text{target}})_{k}$. We train the model using a cross-entropy loss and the Adam optimizer [55], with a batch size of 256. An overview of this process is shown in Figure 2.
Figure 2: In Stage 1, an MLP classifier selects the most suitable circuit topology from a library of human-designed netlists, conditioned on the target performance specification.
Table 1: Classification performance on topology selection.
Metric Score (%)
Accuracy 99.57
Balanced Accuracy 99.33
Macro Precision 99.27
Macro Recall 99.33
Macro F1 99.30
Micro F1 99.57
Evaluation. We begin by assessing the quality of the input representation used for topology classification. Normalized performance vectors encode rich semantic information about circuit behavior. To validate this, we project them into a two-dimensional t-SNE space [56] (Figure 3(a)). The resulting clusters align closely with topology labels, indicating that performance specifications reflect underlying schematic structure and are effective inputs for supervised classification.
We assess classification performance using accuracy, balanced accuracy, macro precision, macro recall, macro F1, and micro F1 scores on the test set. As summarized in Table 1, the classifier achieves an overall accuracy of 99.57%, with macro F1 of 99.30% and balanced accuracy of 99.33%, demonstrating strong generalization across all 20 circuit topologies. Micro F1 (identical to accuracy in the multiclass setting) reaches 99.57%, while macro metrics—averaged equally across classes—highlight robustness to class imbalance. These trends are reinforced by the per-class accuracy plot in Figure 3(c), where most topologies reach 100% accuracy.
The confusion matrix in Figure 3(b) visualizes only the misclassified instances, as most classes achieve perfect accuracy. The few observed errors are primarily concentrated among the two voltage amplifier topologies—common-gate (CGVA) and common-source (CSVA). These circuits operate near the gain-bandwidth limit of the transistor, and when the main amplifier transistor size is held constant, performance metrics such as power consumption, gain, and bandwidth can converge across these architectures. This occasional overlap in the performance space introduces ambiguity in classification for a small subset of instances. For other circuit categories, no significant confusion is expected or observed. These results validate our hypothesis that performance vectors contain sufficient semantic structure for accurate, scalable topology classification.
(a) t-SNE of performance vectors
(b) Confusion matrix (errors only)
(c) Per-class accuracy across circuit topologies
Figure 3: Topology selection results. (a) Performance vectors form well-separated clusters in t-SNE space, showing that circuit functionality is semantically predictive of topology. (b) Misclassifications primarily occur among voltage amplifier variants with overlapping gain-bandwidth tradeoffs. (c) Per-class test accuracy exceeds 93% across all 20 circuit topologies. [^3] The 20 circuit topologies—listed in the same order as the numerical labels in Figure 3(a)—are: CGLNA (Common Gate), CLNA (Cascode), CSLNA (Common Source), DLNA (Differential), DBAMixer (Double-Balanced Active), DBPMixer (Double-Balanced Passive), SBAMixer (Single-Balanced Active), SBPMixer (Single-Balanced Passive), ClassBPA (Class-B), ClassEPA (Class-E), DohPA (Doherty), DPA (Differential), CGVA (Common Gate), CSVA (Common Source), CVA (Cascode), SFVA (Source Follower), IFVCO (Inductive-Feedback), CCVCO (Cross-Coupled), ColVCO (Colpitts), RVCO (Ring).
5 Stage 2: Generalizable Forward Modeling for Performance Prediction
Task Setup. The goal of Stage 2 is to learn a differentiable approximation of the circuit simulator that maps a topology $T$ and parameter vector $x$ to a performance prediction $\hat{y}=f_{\theta}(T,x)$, where $\hat{y}\in\mathbb{R}^{16}$. Unlike black-box simulators, this learned forward model enables efficient performance estimation and supports gradient-based parameter inference in Stage 3. The model is trained to generalize across circuit families and can be reused on unseen topologies with minimal fine-tuning.
Each datapoint consists of a graph-structured Cadence netlist annotated with resolved parameter values and the corresponding performance metrics. We frame the learning task as a supervised regression problem. Since not all performance metrics apply to every topology (e.g., oscillation frequency is undefined for amplifiers), we train the model using a masked mean squared error loss:
$\mathcal{L}_{\text{masked}}=\frac{1}{\sum_{i}m_{i}}\sum_{i=1}^{d}m_{i}\cdot(\hat{y}_{i}-y_{i})^{2},$ (2)
where $m_{i}=1$ if the $i$-th metric is defined for the current sample, and zero otherwise.
Model Architecture and Training. Each circuit is represented as an undirected multi-edge graph with voltage nets as nodes and circuit components as edges. All circuit parameters—both fixed and sweepable—are assigned to edges, along with categorical device types and one-hot encoded indicators. For each edge $(u,v)$, these attributes are concatenated to form a unified feature vector $x_{uv}$. The feature set is consistent within each component type but varies across types (e.g., NMOS vs. inductor), reflecting the structure defined in Section 3.2.
Figure 4: In Stage 2, a custom edge-centric GNN maps an undirected multi-edge graph constructed from the circuit netlist to a performance vector.
To account for component heterogeneity, we apply type-specific MLP encoders $\phi_{\text{enc}}^{(t)}$ to each edge feature vector, producing initial embeddings $e_{uv}^{(0)}=\phi_{\text{enc}}^{(t)}(x_{uv})$, where $t$ is the component type. These embeddings are updated via a 4-layer edge-centric message-passing GNN with shared weights. At each layer $\ell$, for each node $u$, we first compute the node hidden state using the edge embeddings of all neighbors of the node $u$, $\mathcal{N}(u)$. Then, for each edge $(u,v)$ in the circuit graph, we compute the edge embedding at the next layer $\ell+1$ is using the edge embedding $e_{uv}^{(\ell)}$ and the hidden node states forming the edge $(u,v)$ at the current layer $\ell$ as follows:
$h_{u}^{(\ell)}=\sum_{w\in\mathcal{N}(u)}\phi_{\text{MSG}}(e_{wu}^{(\ell)}),\quad e_{uv}^{(\ell+1)}=\phi_{\text{UPD}}(e_{uv}^{(\ell)},h_{u}^{(\ell)},h_{v}^{(\ell)}),$
where $\phi_{\text{MSG}},\phi_{\text{UPD}}$ are the message and update parameters of message-passing GNN and $h_{u}^{(\ell)},h_{v}^{(\ell)}$ are the hidden states for the nodes forming the edge $(u,v)$ respectively. After message passing [57], final edge embeddings $e_{uv}^{(L)}$ are aggregated to form a graph-level representation $z_{\text{graph}}=\sum_{(u,v)}e_{uv}^{(L)}$, which is decoded by a fully connected MLP (hidden size 256) to predict the 16-dimensional performance vector $\hat{y}\in\mathbb{R}^{16}$. An overview of this GNN-based forward prediction pipeline is shown in Figure 4.
To stabilize training, physical parameters are rescaled by their expected units (e.g. resistance by $10^{3}$), and performance targets are normalized to z-scores using training statistics. We train the model using the Adam optimizer (learning rate $10^{-3}$, batch size 256) and a ReduceLROnPlateau scheduler. Xavier uniform initialization is used for all layers, and early stopping is based on validation loss. We adopt the same splits as in Section 4 for consistency in evaluation.
Evaluation. We evaluate the accuracy of the GNN forward model $f_{\theta}$ on a test set drawn from 19 of the 20 topologies. One topology—RVCO—is entirely excluded from training, validation, and test splits to assess generalization to unseen architectures. Prediction quality is measured using standard regression metrics: coefficient of determination ($\text{R}^{2}$), root mean squared error (RMSE), and mean absolute error (MAE), computed independently for each of the 16 performance metrics. We also report the mean relative error per metric, computed as the average across all test samples where each metric is defined. As summarized in Table 2, the model achieves high accuracy across all dimensions, with an average $\text{R}^{2}$ of 0.971.
Figure 5: Distribution of relative error (%) across the test set for the GNN forward model. Plot is trimmed at the 95th percentile.
To evaluate end-to-end prediction accuracy at the sample level, we compute the mean relative error per instance, defined as the average relative error across all valid (non-masked) performance metrics for each test sample. Figure 5 shows the distribution of this quantity across the test set (trimmed at the 95th percentile to reduce the impact of outliers). The distribution is sharply concentrated, indicating that most predictions closely match their corresponding target vectors. Without percentile trimming, the overall mean relative error across the full test set is 9.14%.
Table 2: Prediction accuracy of the forward GNN on all 16 circuit performance metrics.
Metric DCP VGain PGain CGain $\text{S}_{11}$ $\text{S}_{22}$ NF BW OscF TR OutP $\text{P}_{\text{SAT}}$ DE PAE PN VSwg
Unit mW dB dB dB dB dB dB GHz GHz GHz dBm dBm % % dBc/Hz mV
1.0 1.0 0.99 1.0 0.92 1.0 0.99 0.98 0.97 0.83 0.97 1.0 1.0 1.0 0.89 1.0
RMSE 0.289 0.107 0.536 0.84 1.517 0.206 0.534 0.969 0.721 0.293 0.908 0.1 0.232 0.144 2.541 0.07
MAE 0.212 0.077 0.208 0.188 0.554 0.116 0.202 0.369 0.181 0.097 0.232 0.069 0.168 0.104 1.167 0.046
Rel. Err. 11.5% 2.7% 18.6% 7.8% 11.4% 1.8% 4.5% 5.6% 0.6% 6.5% 4.4% 4.5% 4.5% 11.7% 1.3% 1.48%
6 Stage 3: Layout-Aware Parameter Inference via Gradient Reasoning
Task Setup. Given a target performance vector $y_{\text{target}}$ and a selected topology $T^{*}$, the goal of Stage 3 is to recover a parameter vector $x^{*}$ that minimizes a total loss combining performance error and layout-aware penalties, using the learned forward model $f_{\theta}$ from Stage 2. This formulation enables instance-wise inverse design without requiring circuit-level simulation.
To initialize optimization, we perturb domain-specific scale factors (e.g., $10^{-12}$ for capacitors) to sample a plausible starting point $x_{0}$. Parameters are iteratively updated via gradient descent, guided by both functional and physical objectives. Topology-specific constants are held fixed, and parameter values are clipped to remain within valid domain bounds throughout the process.
Loss Function. The total loss follows the structure defined in Eqn 1, jointly minimizing performance mismatch and layout cost:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{perf}}+\lambda_{\text{area}}\cdot\mathcal{L}_{\text{layout}}\cdot g(\mathcal{L}_{\text{perf}}),$ (3)
where $\mathcal{L}_{\text{perf}}$ is the masked mean squared error (see Eqn 2) between predicted and target performance vectors, and $\mathcal{L}_{\text{layout}}$ is a normalized area penalty derived from analytical layout equations. To prioritize functionality, layout loss is softly gated by:
$g(\mathcal{L}_{\text{perf}})=1-\sigma\left(\gamma(\mathcal{L}_{\text{perf}}-\tau)\right),$
which attenuates layout penalties when performance error exceeds a threshold $\tau$, encouraging the model to first achieve functionality before optimizing for layout compactness.
We set $\tau=0.05$, $\gamma=50$, and normalize layout area by $1\,\text{mm}^{2}$ to stabilize gradients. The layout weight $\lambda_{\text{area}}=0.02$ is chosen empirically to balance performance accuracy and physical realism without dominating the loss. This gated formulation supports manufacturable parameter recovery and reflects the broader paradigm of physics-informed learning [58].
Figure 6: In Stage 3, gradient reasoning iteratively updates parameters to minimize a loss combining performance error and layout cost, computed via a differentiable analytical model.
Differentiable Layout Modeling. In mm-wave analog design, layout is not a downstream concern but a critical determinant of circuit performance—particularly for passive components. Substrate coupling, proximity effects, and DRC-imposed geometries directly affect key metrics such as resonance frequency, quality factor, and impedance matching. To incorporate these effects, we introduce a differentiable layout model that computes total physical area analytically from circuit parameters. This enables layout constraints to directly guide parameter optimization during inverse design. By minimizing the layout area in distributed mm-wave circuits [59], unwanted signal loss [60] is reduced, the self-resonance frequency of passives can increase [61], and phase and amplitude mismatches across signal paths [62] can be reduced.
The layout model is deterministic and non-learned. It estimates area contributions from passive components—capacitors, inductors, and resistors—as these dominate total area and exhibit layout-sensitive behavior. Active devices (e.g., MOSFETs) are excluded since their geometries are fixed by the PDK and are negligible [63]. For a given parameter vector $x$, the total layout loss is computed as:
$\mathcal{L}_{\text{layout}}(x)=\sum_{e\in\mathcal{E}_{\text{passive}}}A_{e}(x),$
where $\mathcal{E}_{\text{passive}}$ is the set of passive elements, and $A_{e}(x)$ is the area of the created layout for the passive component based on analytical physics-based equations. The area of element $e$ is estimated based on its 2D dimensions (e.g., $A=W\cdot L$ for resistors and capacitors). This area is normalized and used as a differentiable penalty in the optimization objective (see Eqn 3).
Gradient Reasoning Procedure. Starting from the initialized parameter vector $x$, we iteratively update parameters via gradient reasoning. At each step, the frozen forward model $f_{\theta}$ predicts the performance $\hat{y}=f_{\theta}(T,x)$, and the total loss $\mathcal{L}_{\text{total}}$ is evaluated. Gradients are backpropagated with respect to $x$, and updates are applied using the Adam optimizer. Optimization proceeds for a fixed number of steps, with early stopping triggered if the loss fails to improve over a predefined window.
To handle varying circuit difficulty and initialization quality, we employ an adaptive learning rate strategy. Each instance begins with a moderate learning rate ($10^{-6}$), refined during optimization via a ReduceLROnPlateau scheduler. If the solution fails to meet thresholds on performance error or layout area, optimization restarts with a more exploratory learning rate. This adjustment balances exploration and fine-tuning, enabling rapid convergence to physically valid solutions, typically within milliseconds to under one second per instance. An overview is shown in Figure 6.
Evaluation. We evaluate Stage 3 on 9,500 test instances (500 per topology) using our gradient-based optimization pipeline. A design is considered converged if it meets both: (i) a predicted mean relative error below 10%, and (ii) a layout area under a topology-specific bound—1 mm2 for most circuits and 1.5 mm2 for DLNA, DohPA, and ClassBPA. The 10% error threshold reflects the forward model’s $\sim 9\%$ average prediction error (Section 5). A design is deemed successful if its final Cadence-simulated performance deviates from the target by less than 20%, confirming real-world viability. Our method achieves a success rate of 78.5% and a mean relative error of 17.7% across converged designs, with average inference time under 1 second on a MacBook CPU. Notably, success rate is coupled with the convergence threshold: tighter error bounds yield higher accuracy but require more iterations—critical for large-scale design tasks.
To illustrate the effectiveness of our pipeline, Figure 7 shows a representative result for the DohPA topology: the synthesized schematic is shown on the left, and the corresponding layout is on the right. These results confirm that the recovered parameters are both functionally accurate and physically realizable. Together, they demonstrate that FALCON enables layout-aware inverse design within a single differentiable pipeline—a capability not supported by existing analog design frameworks.
(a) Designed DohPA schematic
(b) Layout of designed DohPA
Figure 7: Stage 3 results for a synthesized DohPA. The schematic (a) reflects optimized parameters to meet the target specification. The layout (b) is DRC-compliant and physically realizable. The final design achieves a mean relative error of 5.4% compared to the target performance.
7 Conclusion and Future Work
We presented FALCON, a modular framework for end-to-end analog and RF circuit design that unifies topology selection, performance prediction, and layout-aware parameter optimization. Trained on over one million Cadence-simulated mm-wave circuits, FALCON combines a lightweight MLP, a generalizable GNN, and differentiable gradient reasoning to synthesize circuits from specification to layout-constrained parameters. FALCON achieves >99% topology selection accuracy, <10% prediction error, and efficient inverse design—all within sub-second inference. The GNN forward model generalizes to unseen topologies with minimal fine-tuning, supporting broad practical deployment.
In future work, we aim to expand the topology library and support hierarchical macroblocks for scalable design beyond the cell level. We also plan to extend the layout-aware optimization with learned parasitic models and EM-informed constraints for more accurate post-layout estimation. Finally, integrating reinforcement learning or diffusion-based models for generative topology synthesis represents a promising step toward general-purpose analog design automation.
Acknowledgments
We thank Andrea Villasenor and Tanqin He for their assistance with circuit data generation. We also thank Mohammad Shahab Sepehri for his insightful discussions and thoughtful feedback during the development of this work.
Appendix A Qualitative Comparison with Prior Work
To contextualize FALCON within the broader landscape of analog circuit design automation, we provide a qualitative comparison against representative prior works in Table 3. This comparison spans key capabilities including topology selection, parameter inference, performance prediction, layout awareness, and simulator fidelity. We additionally assess reproducibility via dataset and code availability, and introduce a new axis—RF/mm-wave support—to highlight methods evaluated on high-frequency circuit blocks such as LNAs, mixers, and VCOs. Compared to existing approaches, FALCON is the only framework that unifies all these dimensions while maintaining foundry-grade fidelity and open-source accessibility. Definitions for each comparison axis are provided in Table 4.
Table 3: Qualitative comparison of FALCON with prior works across key capabilities in analog circuit design automation.
Method Topology Selection Parameter Inference Performance Prediction Layout Awareness Foundry Grade RF/ mm-Wave Public Dataset Public Code
CktGNN [9] ✘ (SPICE)
LaMAGIC [19] ✘ (SPICE)
AnalogCoder [20] ✘ (SPICE)
GCN-RL [10] ✔ (SPICE/Cadence)
Cao et al. [21] ✔ (SPICE/Cadence)
BO-SPGP  [15] ✔ (Cadence)
ESSAB  [22] ✔ (Cadence)
AICircuit [23, 24] ✔ (Cadence)
Krylov et al. [11] ✘ (SPICE)
Deep-GEN  [16] ✘ (SPICE)
Liu et al. [25] ✘ (SPICE + Parasitic Model)
ALIGN [26] ✔ (Cadence)
LayoutCopilot [27] ✔ (Cadence)
AnalogGym [18] ✘ (SPICE)
AutoCkt [13] ✔ (Cadence) ✘ (incomplete)
L2DC  [12] ✔ (Cadence)
CAN-RL  [14] ✔ (Cadence)
AnGeL. [17] ✘ (SPICE)
FALCON (This work) ✔ (Cadence)
Table 4: Definitions of each comparison axis in Table 3.
Column Definition
Topology Selection Does the method automatically select or predict circuit topology given a target specification?
Parameter Inference Does the method infer element-level parameters (e.g., transistor sizes, component values) as part of design generation?
Performance Prediction Can the method predict circuit performance metrics (e.g., gain, bandwidth, noise) from topology and parameters?
Layout Awareness Is layout considered during optimization or training (e.g., via area constraints, parasitics, or layout-informed loss)?
Dataset Fidelity Does the dataset reflect realistic circuit behavior (e.g., SPICE/Cadence simulations, PDK models)?
RF/mm-Wave Is the method evaluated on at least one RF or mm-wave circuit type that reflects high-frequency design challenges?
Public Dataset Is the dataset used in the work publicly released for reproducibility and benchmarking?
Public Code Is the implementation code publicly available and documented for reproducibility?
Appendix B Dataset Details and Performance Metric Definitions
During dataset generation, each simulated circuit instance is annotated with a set of performance metrics that capture its functional characteristics. All simulations are performed at a fixed frequency of 30 GHz, ensuring consistency across circuit types and relevance to mm-wave design. A total of 16 metrics are defined across all circuits—spanning gain, efficiency, impedance matching, noise, and frequency-domain behavior—though the specific metrics used vary by topology. For example, phase noise is only applicable to oscillators. An overview of all performance metrics is provided in Table 5.
B.1 Low-Noise Amplifiers (LNAs)
Low-noise amplifiers (LNAs) are critical components in receiver front-ends, responsible for amplifying weak antenna signals while introducing minimal additional noise. Their performance directly influences downstream blocks such as mixers and analog-to-digital converters (ADCs), ultimately determining system-level fidelity [31]. To capture the architectural diversity of practical radio-frequency (RF) designs, we include four widely used LNA topologies in this study—common-source LNA (CSLNA), common-gate LNA (CGLNA), cascode LNA (CLNA), and differential LNA (DLNA)—as shown in Figure 8.
Table 5: Overview of 16 performance metrics used during dataset generation.
Performance Name Description
DC Power Consumption (DCP) Total power drawn from the DC supply indicating energy consumption of the circuit
Voltage Gain (VGain) Ratio of output voltage amplitude to input voltage amplitude
Power Gain (PGain) Ratio of output power to input power
Conversion Gain (CGain) Ratio of output power at the desired frequency to input power at the original frequency
$\text{S}_{11}$ Input reflection coefficient indicating impedance matching at the input terminal
$\text{S}_{22}$ Output reflection coefficient indicating impedance matching at the output terminal
Noise Figure (NF) Ratio of input signal-to-noise ratio to output signal-to-noise ratio
Bandwidth (BW) Frequency span over which the circuit maintains specified performance characteristics
Oscillation Frequency (OscF) Steady-state frequency at which the oscillator generates a periodic signal
Tuning Range (TR) Range of achievable oscillation frequencies through variation of control voltages
Output Power (OutP) Power delivered to the load
$\text{P}_{\text{SAT}}$ Maximum output power level beyond which gain compression begins to occur
Drain Efficiency (DE) Ratio of RF output power to DC power consumption.
Power-Added Efficiency (PAE) Ratio of the difference between output power and input power to DC power consumption
Phase Noise (PN) Measure of oscillator stability represented in the frequency domain at a specified offset
Voltage Swing (VSwg) Maximum peak voltage level achievable at the output node
The CSLNA is valued for its simplicity and favorable gain–noise trade-off, especially when paired with inductive source degeneration [30]. The CGLNA, often used in ultra-wideband systems, enables broadband input matching but typically suffers from a higher noise figure [32]. The CLNA improves gain–bandwidth product and reverse isolation, making it ideal for high-frequency, high-linearity applications [33]. The DLNA exploits circuit symmetry to enhance linearity and reject common-mode noise, and is commonly found in high-performance RF front-end designs [34]. The design parameters and performance metrics associated with these topologies are summarized in Table 6.
(a) CSLNA
(b) CGLNA
(c) CLNA
(d) DLNA
Figure 8: Schematic diagrams of the four LNA topologies.
Table 6: LNA topologies with parameter sweep ranges, sample sizes, and performance metrics.
Dataset Type Topology (Code) # of Samples Parameter Sweep Range Performance Metrics (Unit)
LNA CGLNA (0) 52k $\text{C}_{\text{1}}$ [100–600] fF DCP (W) PGain (dB) S11 (dB) NF (dB) BW (Hz)
$\text{C}_{\text{2}}$ [50–300] fF
$\text{C}_{\text{b}}$ [250–750] fF
$\text{L}_{\text{d}}$ [80–580] pH
$\text{L}_{\text{s}}$ [0.5–5.5] nH
$\text{W}_{\text{N}}$ [12–23] µm
CLNA (1) 62k $\text{C}_{\text{1}}$, $\text{C}_{\text{2}}$ [50–250] fF
$\text{L}_{\text{d}}$ [140–300] pH
$\text{L}_{\text{g}}$ [0.4–2] nH
$\text{L}_{\text{s}}$ [50–250] pH
$\text{W}_{\text{N1}}$ [3–5] µm
$\text{W}_{\text{N2}}$ [7–9] µm
CSLNA (2) 39k C [100–300] fF
$\text{L}_{\text{g}}$ [4–6] nH
$\text{L}_{\text{s}}$ [100–200] pH
$\text{W}_{\text{N}}$ [2.5–4] µm
$\text{V}_{\text{gs}}$ [0.5–0.9] V
DLNA (3) 92k $\text{C}_{\text{1}}$ [100–190] fF
$\text{C}_{\text{2}}$ [130–220] fF
$\text{L}_{\text{d}}$ [100–250] pH
$\text{L}_{\text{g}}$ [600–900] pH
$\text{L}_{\text{s}}$ [50–80] pH
$\text{W}_{\text{N1}}$ [4–9.4] µm
$\text{W}_{\text{N2}}$ [5–14] µm
B.2 Mixers
Mixers are fundamental nonlinear components in RF systems, responsible for frequency translation by combining two input signals to produce outputs at the sum and difference of their frequencies. This functionality is essential for transferring signals across frequency domains and is widely used in both transmission and reception paths [35]. To capture diverse mixer architectures, we implement four representative topologies in this work—double-balanced active mixer (DBAMixer), double-balanced passive mixer (DBPMixer), single-balanced active mixer (SBAMixer), and single-balanced passive mixer (SBPMixer)—as shown in Figure 9.
The DBAMixer integrates amplification and differential switching to achieve conversion gain and high port-to-port isolation. Despite its elevated power consumption and design complexity, it is well suited for systems requiring robust performance over varying conditions [36]. The DBPMixer features a fully differential structure that suppresses signal leakage and improves isolation, at the cost of signal loss and a strong local oscillator drive requirement [37]. The SBAMixer includes an amplification stage preceding the switching core to enhance signal strength and reduce noise, offering a balanced performance trade-off with increased power consumption and limited spurious rejection [30]. The SBPMixer employs a minimalist switching structure to perform frequency translation without active gain, enabling low power operation in applications with relaxed performance demands [38]. The parameters and performance metrics for these mixer topologies are listed in Table 7.
(a) DPAMixer
(b) DBPMixer
(c) SBAMixer
(d) SBPMixer
Figure 9: Schematic diagrams of the four Mixer topologies.
Table 7: Mixer topologies with parameter sweep ranges, sample sizes, and performance metrics.
Dataset Type Topology (Code) # of Samples Parameter Sweep Range Performance Metrics (Unit)
Mixer DBAMixer (4) 42k C [1–10] pF DCP (W) CGain (dB) NF (dB) VSwg (V)
[…]
$\text{W}_{\text{N1}}$ [10–30] µm
$\text{W}_{\text{N2}}$ [5–25] µm
DBPMixer (5) 42k C [100–500] fF
R [100–600] $\Omega$
$\text{W}_{\text{N}}$ [10–30] µm
SBAMixer (6) 52k C [1–15] pF
[…]
$\text{W}_{\text{N1}}$ [10–30] µm
$\text{W}_{\text{N2}}$ [10–20] µm
$\text{I}_{\text{tail}}$ [3–10] mA
SBPMixer (7) 44k C [1–30] pF
[…]
$\text{W}_{\text{N}}$ [5–29.5] µm
B.3 Power Amplifiers (PAs)
Power amplifiers (PAs) are the most power-intensive components in radio-frequency (RF) systems and serve as the final interface between transceiver electronics and the antenna. Given their widespread use and the stringent demands of modern communication standards, PA design requires careful trade-offs across key performance metrics [39]. Based on the transistor operating mode, PAs are typically grouped into several canonical classes [40]. In this work, we implement four representative topologies—Class-B PA (ClassBPA), Class-E PA (ClassEPA), Doherty PA (DohPA), and differential PA (DPA)—as shown in Figure 10.
The ClassBPA employs complementary transistors to deliver high gain with moderate efficiency, making it suitable for linear amplification scenarios [41]. The ClassEPA uses a single transistor configured as a switch, paired with a matching network. By minimizing the overlap between drain voltage and current, this topology enables high-efficiency operation and improved robustness to component variation [30]. The DohPA combines main and peaking amplifiers using symmetric two-stack transistors, maintaining consistent gain and efficiency under varying power levels [42]. The DPA features a two-stage cascode structure designed to maximize gain and linearity, offering a favorable trade-off between output power and power consumption [43]. For this topology, we replace the transformer with a T-equivalent network to simplify modeling and training of the graph neural network. Parameter sweeps and performance metrics for these PAs are listed in Table 8.
(a) ClassBPA
(b) ClassEPA
(c) DohPA
(d) DPA
Figure 10: Schematic diagrams of the four PA topologies.
Table 8: PA topologies with parameter sweep ranges, sample sizes, and performance metrics.
Dataset Type Topology (Code) # of Samples Parameter Sweep Range Performance Metrics (Unit)
PA ClassBPA (8) 35k C [55–205] fF DCP (W) PGain (dB) $\text{S}_{\text{11}}$ (dB) $\text{S}_{\text{22}}$ (dB) $\text{P}_{\text{SAT}}$ (dBm) DE (%) PAE (%)
$\text{L}_{\text{1}}$ [1–1.4] nH
$\text{L}_{\text{2}}$ [1–8.5] pH
[…]
$\text{W}_{\text{N}}$ [10–20] µm
$\text{W}_{\text{P}}$ [3–8] µm
ClassEPA (9) 46k $\text{C}_{\text{1}}$ [100–200] fF
$\text{C}_{\text{2}}$ [500–700] fF
$\text{L}_{\text{1}}$ [100–300] pH
$\text{L}_{\text{2}}$ [100–150] pH
$\text{W}_{\text{N}}$ [15–30] µm
DohPA (10) 120k $\text{C}_{\text{1}}$ [2–3] pF
$\text{C}_{\text{2}}$ [200–300] fF
$\text{C}_{\text{3}}$, $\text{C}_{\text{5}}$ [100–200] fF
$\text{C}_{\text{4}}$ [300–400] fF
$\text{L}_{\text{1}}$, $\text{L}_{\text{5}}$ [100–200] pH
$\text{L}_{\text{2}}$ [350–450] pH
$\text{L}_{\text{3}}$ [500–600] pH
$\text{L}_{\text{4}}$ [150–250] pH
$\text{L}_{\text{6}}$ [300–400] pH
$\text{W}_{\text{N1}}$, $\text{W}_{\text{N2}}$ [6–13] µm
DPA (11) 80k $\text{L}_{\text{ip}}$ [100–500] pH
$\text{L}_{\text{is}}$ [300–700] pH
$\text{L}_{\text{op}}$ [0.8–1.2] nH
$\text{L}_{\text{os}}$ [400–800] pH
$\text{L}_{\text{m}}$ [50–250] pH
$\text{W}_{\text{N1}}$ [6–31] µm
$\text{W}_{\text{N2}}$ [10–35] µm
B.4 Voltage Amplifiers (VAs)
Voltage amplifiers (VAs) are fundamental components in analog circuit design, responsible for increasing signal amplitude while preserving waveform integrity. Effective VA design requires balancing key performance metrics tailored to both RF and baseband operating conditions [44]. In this work, we implement four widely used VA topologies—common-source VA (CSVA), common-gate VA (CGVA), cascode VA (CVA), and source follower VA (SFVA)—as shown in Figure 11.
The CSVA remains the most widely adopted configuration due to its structural simplicity and high voltage gain. It is frequently used as the first gain stage in various analog systems [45]. The CGVA is suitable for applications requiring low input impedance and wide bandwidth, such as impedance transformation or broadband input matching [46]. The CVA, which cascades a common-source stage with a common-gate transistor, improves the gain–bandwidth product and enhances stability, making it ideal for applications demanding wide dynamic range and robust gain control [47]. The SFVA, also known as a common-drain amplifier, provides near-unity voltage gain and low output impedance, making it well suited for interstage buffering, load driving, and impedance bridging [48]. Parameter ranges and performance specifications for these VA topologies are listed in Table 9.
(a) CSVA
(b) CGVA
(c) CVA
(d) SFVA
Figure 11: Schematic diagrams of the four VA topologies.
Table 9: VA topologies with parameter sweep ranges, sample sizes, and performance metrics.
Dataset Type Topology (Code) # of Samples Parameter Sweep Range Performance Metrics (Unit)
VA CGVA (12) 33k C [0.1–1.5] pF DCP (W) VGain (dB) BW (Hz)
[…]
$\text{W}_{\text{N1}}$ [5–30] µm
$\text{W}_{\text{N2}}$ [5–10] µm
[…]
$\text{W}_{\text{N}}$ [3–15] µm
$\text{V}_{\text{DD}}$ [1–1.8] V
$\text{V}_{\text{gate}}$ [0.6–0.9] V
[…]
$\text{W}_{\text{N1}}$, $\text{W}_{\text{N2}}$ [1–10] µm
$\text{W}_{\text{N3}}$ [10–15] µm
SFVA (15) 28k $\text{W}_{\text{N1}}$ [40–60] µm
$\text{W}_{\text{N2}}$ [2–8] µm
$\text{V}_{\text{DD}}$ [1.1–1.8] V
$\text{V}_{\text{gate}}$ [0.6–1.2] V
$\text{V}_{\text{b}}$ [0.5–0.9] V
B.5 Voltage-Controlled Oscillators (VCOs)
Voltage-controlled oscillators (VCOs) are essential building blocks in analog and RF systems, responsible for generating periodic waveforms with frequencies modulated by a control voltage. These circuits rely on amplification, feedback, and resonance to sustain stable oscillations. Owing to their wide tuning range, low power consumption, and ease of integration, VCOs are broadly used in systems such as phase-locked loops (PLLs), frequency synthesizers, and clock recovery circuits [49]. In this work, we implement four representative VCO topologies—inductive-feedback VCO (IFVCO), cross-coupled VCO (CCVCO), Colpitts VCO (ColVCO), and ring VCO (RVCO)—as shown in Figure 12.
The IFVCO employs an NMOS differential pair with an inductor-based feedback path to sustain oscillations. This topology provides favorable noise performance and compact layout, making it well suited for low-voltage, low-power designs [50]. The CCVCO achieves negative resistance through cross-coupling, enabling low phase noise and high integration density, and is widely adopted in frequency synthesizers and PLLs [51]. The ColVCO uses an LC tank and capacitive feedback to achieve high frequency stability and low phase noise, making it ideal for precision RF communication and instrumentation [52]. The RVCO consists of cascaded delay stages forming a feedback loop, offering low power consumption, wide tuning range, and minimal area footprint, though at the cost of higher phase noise. It is commonly used in on-chip clock generation and low-power sensor applications [53]. Design parameters and performance metrics for these VCO topologies are presented in Table 10.
(a) IFVCO
(b) CCVCO
(c) ColVCO
(d) RVCO
Figure 12: Schematic diagrams of the four VCO topologies.
Table 10: VCO topologies with parameter sweep ranges, sample sizes, and performance metrics.
Dataset Type Topology (Code) # of Samples Parameter Sweep Range Performance Metrics (Unit)
VCO IFVCO (16) 43k $\text{C}_{\text{1}}$ [700–900] fF DCP (W) OscF (Hz) TR (Hz) OutP (dBm) PN (dBc/Hz)
$\text{C}_{\text{2}}$ [50–250] fF
$\text{L}_{\text{1}}$ [400–600] pH
$\text{L}_{\text{2}}$ [500–700] pH
$\text{W}_{\text{N}}$, $\text{W}_{\text{var}}$ [5–9] µm
CCVCO (17) 54k L [200–400] pH
$\text{W}_{\text{N}}$ [10–35] µm
$\text{W}_{\text{var}}$ [5–30] µm
ColVCO (18) 90k C [80–140] fF
L [250–350] pH
$\text{W}_{\text{N}}$ [30–50] µm
$\text{W}_{\text{var}}$ [5–15] µm
$\text{V}_{\text{b}}$ [0.7–1.2] V
$\text{I}_{\text{tail}}$ [5–15] mA
RVCO (19) 46k C [300–700] fF
$\text{L}_{\text{1}}$ [300–500] pH
$\text{L}_{\text{2}}$ [50–250] pH
$\text{W}_{\text{N}}$ [20–40] µm
$\text{W}_{\text{var}}$ [5–25] µm
Appendix C Graph-Based Circuit Representation
To enable GNN-based modeling of analog circuits, we represent each netlist as a directed multigraph where nodes correspond to electrical nets and edges encode circuit components such as transistors, capacitors, inductors, and voltage sources. Each edge is labeled with its component type and terminal role (e.g., gate, source, drain), and component-specific attributes are stored as edge features. For transistors, labels such as GS, DS, and DG denote source-to-gate, drain-to-source, and drain-to-gate connections, respectively.
Figure 13 illustrates two representative graph structures extracted from our dataset: an IFVCO and a ClassBPA. The visual encoding highlights the diversity of components and connectivity patterns across topologies. Edges corresponding to the same component type share a common color for visual consistency and semantic clarity. These structured graphs serve as the primary input to our GNN pipeline for performance prediction and inverse design.
(a) IFVCO
(b) ClassBPA
Figure 13: Graph representations of two analog circuit topologies from our dataset: (a) IFVCO and (b) ClassBPA. Nodes represent electrical nets, and colored edges denote circuit components such as transistors, capacitors, inductors, and sources. Each component type is visually distinguished by color and labeled with its name and terminal role (e.g., N2_GS, V0). These graphs serve as input to our GNN-based performance modeling and inverse design pipeline.
Appendix D Generalizing to Unseen Topologies via Fine-Tuning
To assess the generalization ability of our pretrained GNN, we evaluate it on the held-out RVCO topology, which was entirely excluded from the Stage 2 training, validation, and test splits (see Section 5). Notably, the RVCO training partition used here matches that of the Stage 1 experiments (Section 4), enabling consistent cross-stage evaluation.
We fine-tune the GNN by freezing all encoder and message-passing layers and updating only the final output head (output_mlp). Fine-tuning is performed on the RVCO training set, which contains approximately 30,000 instances, and completes in under 30 minutes on a MacBook CPU.
Even in the zero-shot setting—where the model has never seen RVCO topologies—the pretrained GNN achieves a nontrivial mean relative error of 33.1%, highlighting its strong cross-topology generalization. Fine-tuning reduces this error to just 0.8%, demonstrating that the structural and parametric priors learned during pretraining are highly transferable. Table 11 reports detailed performance across five key metrics, confirming that the pretrained GNN can be rapidly adapted to novel circuit families with minimal supervision.
Table 11: Fine-tuning results on the held-out RVCO topology. Only the output head is updated using RVCO samples.
Metric DCP OscF TR OutP PN
Unit W GHz GHz dBm dBc/Hz
1.0 1.0 1.0 0.97 0.99
RMSE 0.725 0.244 0.022 0.098 0.792
MAE 0.576 0.19 0.018 0.078 0.477
Rel. Err. 0.85% 0.62% 1.4% 0.69% 0.56%
Appendix E Layout Design and DRC Compliance
E.1 Design Rule Enforcement in 45 nm CMOS
We implemented FALCON using a 45 nm CMOS technology node, applying rigorous Design Rule Checking (DRC) at both the cell and full-chip layout levels. At the cell level, our parameterized layout generators enforced foundry-specific constraints, including minimum feature width and length, contact and via spacing, and metal enclosure rules. At the circuit level, we incorporated physical verification to mitigate interconnect coupling, IR drop, and layout-dependent parasitic mismatches—factors that are especially critical in high-frequency and precision analog design.
DRC plays a vital role in ensuring that layouts comply with process design rules defined by the semiconductor foundry. Adhering to these rules ensures not only physical manufacturability but also electrical reliability. Violations may lead to fabrication failures, including yield degradation, electrical shorts or opens, electromigration-induced issues, and parasitic mismatches. Moreover, DRC compliance is essential for compatibility with downstream fabrication steps such as photomask generation, optical lithography, and chemical-mechanical planarization (CMP), safeguarding the yield and fidelity of the final IC.
Circuit-Level Layout Guidelines. We enforced several topology-aware layout constraints during full-circuit integration to preserve signal integrity and robustness:
  • Inductor-to-inductor spacing: $\geq$ 35.0 $\mu$m to mitigate mutual inductive coupling and magnetic interference.
  • Guardring placement: Sensitive analog blocks are enclosed by N-well or deep N-well guardrings with spacing $\geq$ 5.0 $\mu$m to suppress substrate noise coupling.
  • Differential pair symmetry: Differential signal paths are layout-matched to ensure $\Delta L<0.5\,\mu$m, minimizing mismatch and preserving phase balance.
DRC Constraints and Layer Definitions. Table 12 summarizes the DRC constraints applied to key analog components across relevant process layers. Table 13 provides the abbreviations used for metal, contact, and via layers in the 45 nm CMOS process.
Table 12: Design rule constraints for key analog components in 45 nm CMOS.
Component Layer Physical Constraint Symbol Value Unit
MIM Capacitor (QT, LD, VV, OB) QT/LD Minimum Cap Width $W_{\text{MIN}}$ 6.05 $\mu$m
QT/LD Maximum Cap Width $W_{\text{MAX}}$ 150.0 $\mu$m
QT/LD Cap Length $L$ 6.0 $\mu$m
VV VV Square Size VV_SIZE 4.0 $\mu$m
VV VV Spacing VV_SPACE 2.0 $\mu$m
VV VV to Edge Spacing VV_EDGE_MIN 1.0 $\mu$m
Resistor (RX, CA, M1) RX Minimum Width $W_{\text{MIN}}$ 0.462 $\mu$m
RX Maximum Width $W_{\text{MAX}}$ 5.0 $\mu$m
RX Minimum Length $L_{\text{MIN}}$ 0.4 $\mu$m
RX Maximum Length $L_{\text{MAX}}$ 5.0 $\mu$m
CA Contact Size CA_SIZE 0.06 $\mu$m
CA Contact Spacing CA_SPACE 0.10 $\mu$m
CA CA to Edge Spacing CA_EDGE 0.11 $\mu$m
Inductor (M3) M3 Minimum Width M3_W_MIN 2.0 $\mu$m
M3 Maximum Width M3_W_MAX 20.0 $\mu$m
M3 Minimum Spacing M3_S_MIN 2.0 $\mu$m
Grid All Layers Minimum Grid Min_Grid 0.005 $\mu$m
Table 13: Process layer abbreviations in the 45 nm CMOS design flow.
Layer Name Description
RX Resistor implant or diffusion layer used to define integrated resistor geometries.
CA Contact layer forming vias between diffusion/poly and the first metal layer (M1).
M1 First metal layer, typically used for local interconnects and resistor terminals.
M3 Third metal layer, used for wider routing tracks and planar inductor layouts.
QT Top metal plate in MIM capacitor structures, providing the upper electrode.
LD Lower metal plate in MIM capacitor structures, acting as the bottom electrode.
VV Via layer connecting different metal layers, especially in capacitor and dense routing regions.
OB Opening/blocking layer used to define restricted zones, often to exclude metal or for CMP mask clarity.
E.2 MIM Capacitor Capacitance Model
The total capacitance $C_{N}$ of a metal-insulator-metal (MIM) capacitor is modeled as:
[…]
where $L$ and $W$ are the layout length and width in $\mu$m, $C_{a}$ is the area capacitance density, and $C_{p}$ is the fringing field contribution per unit length. This model includes both area and perimeter contributions to more accurately reflect layout-dependent capacitance in IC design (see Figure 14(a)).
(a) MIM capacitor layout
(b) Resistor layout
(c) Inductor layout
Figure 14: Layout views of passive components. (a) MIM capacitor with metal-insulator-metal stack. (b) Resistor layout with matching geometry. (c) Spiral inductor with octagonal turns for optimized area and Q-factor.
1. Area Capacitance Term:$C_{a}\cdot L\cdot W$
Physical Concept: This term represents the primary (parallel-plate) capacitance formed between the overlapping top and bottom metal layers. It arises from the uniform electric field across the dielectric.
Layer Physics Explanation:
  • $L\cdot W$ corresponds to the overlap area of the plates.
  • $C_{a}=0.335\,\text{fF}/\mu\text{m}^{2}$ is the area capacitance density, derived from:
    • Dielectric permittivity $\varepsilon$ of the insulating material.
    • Dielectric thickness $d$, with $C\propto\varepsilon/d$.
2. Perimeter (Fringing) Capacitance Term:$C_{p}\cdot 2\cdot(L+W)$
Physical Concept: This term models fringing fields at the plate edges, contributing additional capacitance—particularly relevant in small geometries.
Layer Physics Explanation:
  • $2\cdot(L+W)$ is the physical perimeter of the capacitor.
  • $C_{p}=0.11\,\text{fF}/\mu\text{m}$ accounts for the fringing field contribution per unit length.
Summary: This composite model enables accurate estimation of MIM capacitance by capturing both parallel-plate and fringing effects. The constants $C_{a}$ and $C_{p}$ are typically calibrated using process-specific measurements or electromagnetic simulations.
[…]
The corresponding bounding area is estimated from the component’s geometric envelope:
[…]
E.3 N+ Silicided Polysilicon Resistor Model
The resistance of a layout-defined resistor implemented using the ndslires layer is modeled as:
$R=R_{s}\cdot\frac{L}{W+\Delta W}+2R_{\text{end}}+\delta\quad[\Omega]$
Physical Concept: This structure uses heavily doped N+ polysilicon overlaid with a silicide layer to reduce resistance. Current flows laterally through the poly-silicide film (see Figure 14(b)), and resistance is shaped by the aspect ratio of the layout as well as process-dependent corrections.
Layer Physics Explanation:
  • $R_{s}=17.6\,\Omega/\square$ (ohm per square) is the sheet resistance of the silicided poly layer.
  • $W=5.0\,\mu$m is the drawn width; $\Delta W=0.048\,\mu$m accounts for process-induced width bias.
  • $L$ is the drawn resistor length.
  • $R_{\text{end}}=1\,\Omega$ models terminal resistance due to contact diffusion and current crowding.
  • $\delta=0.917\,\Omega$ accounts for residual layout-dependent parasitics.
Summary: The empirical layout relation used in parameterized generation is:
$\boxed{R\approx 3.5007\cdot L+2.917\quad[\Omega]}$ (6)
[…]
E.4 Octagon Spiral Inductor Model
Physical Concept: Accurate modeling and layout optimization of planar spiral inductors are critical in analog circuit design. Inductor performance is highly sensitive to parasitic elements, achievable quality factor ($Q$), and layout constraints imposed by process design rules. To support accurate performance prediction and inform layout choices, we adopt a modified power-law model that expresses inductance as a function of key geometric parameters. The model is validated against empirical measurements and shows strong agreement with classical analytical formulations.
Numerous classical formulations relate inductance to geometric factors such as the number of turns, average diameter, trace width, and inter-turn spacing. Among these, the compact closed-form expressions in RF Microelectronics textbook [30] are widely adopted for their balance of simplicity and accuracy. Building on this foundation, we adopt a reparameterized monomial model that better fits our empirical measurement data:
$L=2.454\times 10^{-4}\cdot D_{\text{out}}^{-1.21}\cdot W^{-0.163}\cdot D_{\text{avg}}^{2.836}\cdot S^{-0.049}\quad[$\mathrm{nH}$]$
Layer Physics Explanation:
  • $D_{\text{out}}=2(R+\frac{W}{2})$ is the outer diameter,
  • $D_{\text{in}}=2(R-\frac{W}{2})$ is the inner diameter,
  • $D_{\text{avg}}=(D_{\text{out}}+D_{\text{in}})/2=2R$ is the average diameter,
[…]
L ($\mathrm{nH}$) 0.123 0.170 0.220 0.276
[…]
Summary: With $W$ and $S$ fixed, inductance simplifies to:
$\boxed{L\approx 2.337\times 10^{-3}\cdot R^{1.164}\quad\text{[$\mathrm{nH}$]}}$ (8)
The bounding area is estimated by:
[…]
The performance of on-chip inductors is fundamentally influenced by layout-dependent factors such as trace width, metal thickness, and inter-turn spacing. Increasing the trace width ($W_{\text{ind}}$) reduces series resistance by enlarging the conductor’s cross-sectional area, thereby improving the quality factor, $Q=\omega L/R_{\text{series}}$. However, wider traces also increase parasitic capacitance to adjacent turns and the substrate, which lowers the self-resonance frequency.
[…]
Turn-to-turn spacing ($S$) affects both inductance and quality factor ($Q$). Tighter spacing enhances magnetic coupling, thereby increasing inductance density. However, it also intensifies capacitive coupling and dielectric losses—particularly in modern CMOS processes with high-$k$ inter-metal dielectrics—which can degrade $Q$. Conversely, excessive spacing reduces inductance without providing a proportionate benefit in loss reduction. As a result, one-turn spiral inductors are commonly favored in RF design due to their low series resistance, minimized parasitics, and improved modeling predictability.
[…]
The implemented spiral inductor geometry is shown in Figure 14(c). Table 15 summarizes the DRC-compliant tuning ranges, estimated layout areas, and decomposition strategies for single-cell passive components in our layout library.
Table 15: Single-cell passive component limits based on DRC and associated layout area costs.
Component Tunable Variable Value Range Area Range Decomposition Rule
Resistor Length $L$ 4.32–20.42 $\Omega$ 10.44–34.36 $\mu\mathrm{m}^{2}$ Series if $>$ max, parallel if $<$ min
Capacitor Width $W$ 46.32–1042.4 fF 176–3344 $\mu\mathrm{m}^{2}$ Parallel if $>$ max, series if $<$ min
Inductor Radius $R$geq 0.1 nH $\geq$ 5640 $\mu\mathrm{m}^{2}$ Continuous radius scaling
E.5 Layout Examples of Synthesized Circuits
To illustrate the correspondence between schematic and layout representations, we present two representative synthesized circuits: an IFVCO and a DLNA, shown in Figure 15 and Figure 16, respectively.
In the IFVCO example, the inductor labeled $L_{3}$ functions as an RF choke and is excluded from the on-chip layout due to its large area requirement. Instead, it is intended for off-chip implementation at the PCB level and connected to the die via wire bonding. This external connection is indicated by the yellow pad in Figure 15(b), which serves as the wire-bonding interface.
Since the current stage of system lacks automated routing, all interconnects in the layout were manually drawn to ensure accurate correspondence with the schematic connectivity. These examples demonstrate that synthesized circuit parameters can be successfully translated into DRC-compliant, physically realizable layouts, bridging the gap between high-level optimization and tapeout-ready design.
(a) Designed IFVCO schematic
(b) Layout of designed IFVCO
Figure 15: Stage 3 results for a synthesized IFVCO. The schematic (a) reflects optimized parameters to meet the target specification. The layout (b) is DRC-compliant and physically realizable. The final design achieves a mean relative error of 1.3% compared to the target performance.
(a) Designed DLNA schematic
(b) Layout of designed DLNA
Figure 16: Stage 3 results for a synthesized DLNA. The schematic (a) reflects optimized parameters to meet the target specification. The layout (b) is DRC-compliant and physically realizable. The final design achieves a mean relative error of 5.0% compared to the target performance.
Appendix F Practical Considerations and Limitations
F.1 Training and Inference Efficiency
Although our codebase supports GPU acceleration, all experiments in this work—excluding initial dataset generation—were conducted entirely on a MacBook CPU. This highlights the efficiency and accessibility of the FALCON pipeline, which can be executed on modest hardware without specialized infrastructure. Our MLP and GNN models contain 207k and 1.4M trainable parameters, respectively, with memory footprints of just 831 KB and 5.6 MB.
In Stage 1, the MLP classifier trains in approximately 30 minutes with a batch size of 256 and performs inference in the order of milliseconds per batch. Stage 2’s GNN model takes around 3 days to train on the full dataset using the same batch size and hardware. Fine-tuning on an unseen topology (e.g., RVCO) using $\sim$30,000 samples completes in under 30 minutes.
In Stage 3, the pretrained GNN is used without retraining to perform layout-constrained parameter inference via gradient-based optimization. Inference is conducted one instance at a time (batch size 1), with typical runtimes under 1 second per circuit. Runtime varies based on the convergence threshold and circuit complexity but remains below 2–3 seconds in the worst case across the full test set.
A solution is considered successful if the predicted performance meets the target within a specified relative error threshold. While tighter thresholds (e.g., 5%) improve accuracy, they require more optimization steps—particularly over large datasets. As a result, both success rate and inference time in Stage 3 are directly influenced by this tolerance, which can be tuned based on design fidelity requirements.
F.2 Limitations
This work focuses on a representative set of 20 curated analog topologies spanning five circuit families. While this enables rigorous evaluation and benchmarking, extending support to arbitrary user-defined netlists is a natural next step. Notably, our GNN demonstrates strong generalization even to topologies not seen during training (e.g., RVCO), suggesting broad applicability within this space.
All interconnect routing in the layout phase is currently performed manually, as the flow does not yet integrate a full analog router. This decision allows precise control over parasitic management and DRC compliance but limits scalability for more complex designs. Integration with automated layout tools can further streamline this process.
We do not perform repeated training runs for each model configuration, as the pipeline is designed to be computationally efficient and executable on CPU-only hardware. Unlike benchmarking-focused pipelines, our flow incorporates analog simulation and layout-aware inference, both of which are costly to rerun at scale. Moreover, simulation and layout processes are deterministic, eliminating the need for result averaging over multiple random seeds.